Fay throw a basketball from the basketball hoop. The quadratic equation that models the path of the ball is p(t)= -16t(to the 2nd power) + 20t + 6. If the is 10 feet high, how long is the ball in the air before the ball goes through the hoop?
Answer by stanbon(75887)

Fay throw a basketball from the basketball hoop.
The quadratic equation that models the path of the ball is p(t)= -16t(to the 2nd power) + 20t + 6.
If the is 10 feet high, how long is the ball in the air before the ball goes through the hoop?
------
Draw the picture.
The ball starts at a height of 6 ft., reaches and peak height, then
descends to the basket at a height of 10 ft.
-----
Solve -16t^2+20t+6 = 10
-16t^2 + 20t-4 = 0
4t^2 - 5t + 1 = 0
(4t-1)(t-1) = 0
t = 1/4 or to = 1
---
The ball is at height of 10 ft after 1/4 second (before it
reaches its peak height) and after 1 second (when it hits the basket).
Answer: 1 second
============================
Cheers,
Stan H.